# references.bib

### v1

@@ 2 Experiments
The benefits of any new model selection method can only be verified when the oracle criterion suggests there is “room for improvement” over a basic random split, i.e., there is a performance gap between the two. Thus, the experiments described in this section are set up to reflect this scenario (the limitations of this are discussed further in Section 2.5). For example, it is noticed that UDA tends to exhibit a larger gap than DG, and this is especially pronounced (perhaps unsurprisingly) for adversarial algorithms, which tend to be more sensitive to hyperparameter choices.
Two batches of experiments are run, to reflect both the UDA and DG settings. Each batch comprises an identical training setup applied to 3 different datasets. For the DG experiments, models are trained using the CORAL algorithm [Sun2016DeepAdaptation], with the clustering performed using constraints (1). For the UDA experiments, the DANN algorithm [Ganin2015Domain-AdversarialNetworks] is used to adapt to an additional, unlabelled subset of test domain samples, as well as to align the training domains to each other. As DANN was observed to be more sensitive to domain imbalances, the validation split is set to preserve domain distributions, i.e., using constraints (2).
All feature extractors are finetuned on the entirety of $S$ before the features are computed for the clustering, regardless of pretraining. Experiments are conducted using the DomainBed framework [Gulrajani2021InGeneralization], with the Gurobi Optimizer [GurobiOptimizationLLC2023GurobiManual] used to solve the LPs. Further training and hyperparameter search details are given in Appendix A. In total, the experiments involve training 5,160 models, requiring around 100 GPU-days of computation.
@@ 2.1 Datasets
2.1.1 DG
Camelyon17-WILDS [Bandi2019FromChallenge, Koh2021WILDS:Shifts] tumour detection in tissue samples across 5 hospitals, 2 classes and 455,954 samples. In keeping with the WILDS setup, the model in this case is trained from scratch rather than using pretrained weights. However, note that these results still cannot be compared directly with results from WILDS, as the DomainBed setup does not match exactly. License: CC0.
Humpbacks [Napoli2023UnsupervisedCalls] detection of humpback whale vocalisations across 4 recording locations, 2 classes and 43,385 samples. This is the only dataset not to use the ResNet-18 architecture, and instead uses a custom CNN architecture and acoustic front-end described in Appendix A. License: Proprietary.
SVIRO [DiasDaCruz2020SVIRO:Benchmark] classification of vehicle rear seat occupancy across 10 car models, 7 classes and a balanced subset of 24,500 samples of the original dataset. License: CC BY-NC-SA 4.0.
2.1.2 UDA
VLCS [Fang2013UnbiasedBias] object classification across 4 image datasets, 5 classes and 10,729 samples. License: unknown.
PACS [Li2017DeeperGeneralization] object classification across 4 image styles (photos, art, cartoons, and sketches), 7 classes and 9,991 samples. License: unknown.
Terra Incognita [Beery2018RecognitionIncognita] classification of wild animals across 4 camera trap locations, 10 classes and 24,788 samples. License: CDLA-Permissive 1.0.
With the exception of Camelyon17, the datasets are all small enough that the entire kernel matrix can be computed. So, Camelyon17 is the only dataset for which the Nyström method is applied.
@@ 2.2 Results
In total, 6 model selection methods are compared. These are: the random split; the leave-one-domain-out split; the test domain (oracle) validation set; a random split followed by mix-up augmentation on the validation set [Lu2023TowardsSolution]; and two variants of the cluster-based split described in Section 1, using a linear kernel $\kappa(x,y)=x^{T}y$ (linear k-means); and a radial basis function (RBF) kernel $\kappa(x,y)=e^{-\gamma\left\|x-y\right\|^{2}}$ with bandwidth parameter $\gamma=1$. The results are shown in Table 1, along with standard errors (the standard errors are as computed by DomainBed, and capture variability in the overall experimental run, including random seeds and across domains). In the following sections, a 95% confidence level is used when verifying whether two values have a statistically significant difference, which corresponds to non-overlapping confidence intervals of 1.96 times the standard error (assuming normally distributed errors).
@@
Split type Camelyon17 Humpbacks SVIRO VLCS PACS TerraInc
@@
The range of possible performance improvement differs by dataset, as determined by the gap between the random split and oracle criterion; accuracy values should be considered relative to this scale when averaging across datasets. Therefore, a column of average normalised values is also shown, where each dataset is shifted and scaled to give the random split a value of 0 and the oracle a value of 100.
@@
Where the validation set comprises multiple domains, model selection is based on average validation accuracy across these domains, as is the DomainBed default. Results based on worst-domain accuracy are also shown in Appendix B. Overall, no significant difference in test accuracy is seen if worst-case validation accuracy is used, although the standard error does increase.
@@
On average, the cluster-based splits provide a net absolute accuracy gain of around 3 percentage points compared to the random split, and 2 percentage points gain compared to leave-one-domain out validation. In relative terms, clustering is observed to close around 50% of the gap between the random split and oracle criterion, compared to 28% for leave-one-domain-out validation and 10% for mix-up. Overall, performance is slightly higher using the linear kernel than with the RBF kernel, although this is within margin of error. It is possible that the latter may be improved by taking more care into choosing a suitable value for $\gamma$.
An ablation study comparing the effects of different clustering parameters, on the VLCS dataset, is conducted in Appendix C.
2.3 MMD analysis
As stated in Section 1, the motivation for cluster-based splits is the assumption that increasing the MMD between the training and validation sets increases test domain accuracy. To provide empirical support for this, these two variables are plotted against each other in Figure 1, for each dataset. Again, the Nyström method is used to estimate the MMD for the Camelyon17 dataset due to its size. Each subplot in Figure 1 shows a different dataset, and each point in a subplot represents one of the 5 split types (not including the oracle). The correlation coefficients and associated $p$-values are also shown. As only monotonic associations are being tested for, Spearman’s rank correlation $\rho$ is used.
Even with only 5 datapoints, Camelyon17 gives by far the clearest indication of a logarithmic-style relationship between the MMD and test accuracy, possibly because the dataset is large enough that a low-noise estimate of the MMD is possible. The results for this dataset are given in tabular form in Appendix D, along with an additional ablation comparing the different clustering constraints (1) and (2).
The results for all datasets plotted on a single set of axes are shown in Figure 2. As the ranges of both the accuracy and MMD values vary across datasets, these must be normalised in order to combine them. The median-IQR method (where each dataset is scaled to have zero median and unit interquartile range, on both axes) is used as this is the most robust to the large outliers mentioned. The MMDs are plotted on a symmetrical log scale [Webber2013AData] to make the correlation more visually evident. The means for each split type are also indicated.
Spearman’s $\rho$ for the combined datasets is 0.63, indicating a strong positive correlation between the MMD and test accuracy, whilst a $p$-value in the order of $10^{-4}$ indicates with strong significance that a rank correlation exists. Furthermore, Pearson’s $r$ coefficient for the log-scaled data in Figure 2 is calculated to be 0.61, with a $p$-value of $3.79\times 10^{-4}$. This $p$-value is for the test of the stronger hypothesis that the relationship between the two normalised variables is symmetrical-logarithmic, and again is strongly indicative that this is the case.
The evidence in this section supports the proposition that the validation split should be attempting to maximise the MMD between the training and validation sets, and that this is more effective than trying to create a domain shift of the same “type” as that of the test set, as a traditional metadata-based OOD validation split may do.
Figure 2: The MMD between the training and validation sets versus test domain accuracy, normalised by dataset, and plotted on a single set of axes. The mean for each split type is indicated with an asterisk (*). Note there are 6 overlaying datapoints at the origin, corresponding to the medians of each dataset.
2.4 Assessing the effect of the changing training distribution
The use of non-random data splitting introduces a confounding variable to the experiments: as both the training and validation distributions are dependent on the split, it is possible the test accuracy is being influenced by the model training, as well as the validation. To support the claim that cluster-based splitting results in more generalisable model selection, it is necessary to decouple the effects of these interventions, and show that it is indeed the improved model validation, and not the training, that is improving test accuracy.
DG experiments UDA experiments
Split type Camelyon17 Humpbacks SVIRO VLCS PACS TerraInc Average
Leave-one
domain out
85.4 ± 0.4 82.1 ± 1.3 99.0 ± 0.1 73.8 ± 0.3 82.2 ± 0.5 38.9 ± 1.2 76.9 ± 0.3
Linear
k-means
88.0 ± 0.3 85.4 ± 0.8 99.1 ± 0.1 77.0 ± 0.3 85.7 ± 0.3 47.1 ± 0.9 80.4 ± 0.2
RBF kernel
k-means
88.6 ± 0.6 85.0 ± 0.7 99.2 ± 0.1 77.9 ± 0.5 85.7 ± 0.3 44.5 ± 0.6 80.2 ± 0.2
These results show that performing a cluster-based split has no significant effect on the generalisation power of a model trained on the training set produced by that split (that is, when the hyperparameters are being chosen via the oracle criterion). Therefore, it can be concluded that the change in performance of this method compared to the random split comes entirely from the model selection. On the other hand, test accuracy for the leave-one-domain-out split does significantly reduce. This finding may help to explain why metadata-based splits have been found to underperform random splits in some works [Gulrajani2021InGeneralization]: the increased robustness of OOD validation is not enough to counterbalance the reduced robustness of training on fewer domains.
2.5 Limitations
As stated in Section 1, a major advantage of clustering-based validation is that it is not dependent on domain metadata. However, even with the domain constraint removed or with metadata-independent training algorithms (e.g., ERM), the specific experimental setup in this paper precludes the ability to test this method in truly metadata-free settings. Due to the fundamental design of the DomainBed framework, domain metadata can still be leveraged (both implicitly and explicitly) through several avenues. For example, models are still trained on domain-balanced minibatches of data, and validation accuracy is averaged over the validation domains. The datasets themselves may also be unrealistically domain-balanced to begin with. However, the authors believe that these influences are mild enough that the overall trends observed in these experiments can reasonably be expected to hold in metadata-free settings as well.
3 Conclusion
The algorithm is not parameter-free; future work could include a data-driven method for selecting these parameters, for example with an additional layer of meta-tuning. The algorithm can also trivially be extended to k-fold cross-validation by increasing the number of clusters.
[…]
Appendix A Additional training and hyperparameter details
Experimental parameter Value
Hyperparameter random search size 10
Number of trials 3
Holdout fraction 0.2
UDA holdout fraction 0.5
Number of training steps 3000
Gaussian kernel bandwidth 1
Finetuning iterations before split 3000
Nyström subset size (if applicable) 2000
Architecture ResNet-18
Class balanced True
Table 4: Classification pipeline of the Humpbacks dataset. This follows the pipeline in [Napoli2023UnsupervisedCalls].
Step number Step detail
Acoustic front-end
1 Resample to 10 kHz
2 Mel-scale filter bank with 64 filters
3 Short-time Fourier transform with 100 ms FFT window, 50% overlap
4 Per-channel energy normalisation
5 Split into 3.92 s (128 pixel) analysis frames with 50% overlap
CNN
1 Conv2D (nodes=16, kernel=3x3, stride=2, activation=ReLU)
2 Conv2D (nodes=16, kernel=3x3, stride=2, activation=ReLU)
3 Conv2D (nodes=16, kernel=3x3, stride=2, activation=ReLU)
4 Conv2D (nodes=16, kernel=3x3, stride=2, activation=ReLU)
5 Global average-pooling 2D
6 Fully-connected layer
Appendix B Worst case accuracy validation
DG experiments UDA experiments
Split type Camelyon17 Humpbacks SVIRO VLCS PACS TerraInc
Raw
average
Normalised
average
Leave-one
domain out
Linear
k-means
RBF kernel
k-means
Appendix C Ablation study
An ablation study conducted on the VLCS dataset is shown in Table 6. This shows the effects of finetuning the feature extractor before clustering, the Nyström approximation, and the different constraint sets (1) and (2).
@@ Appendix C Ablation study
For this dataset, use of the Nyström approximation, as well as additional finetuning of $\theta_{F}$, are not observed to have significant effects on test accuracy. For the linear kernel, clustering with constraints (1) performs significantly lower than using constraints (2), however, for the RBF kernel, this difference is not observed.
Appendix D MMD analysis on Camelyon17
This table shows the MMD versus model accuracy on the Camelyon17 dataset (as in Figure 1), along with an additional ablation comparing the different clustering constraints (1) and (2). It can be seen that the additional constraint on the clustering (i.e., taking $g\in\mathcal{Y\times D}$) reduces the optimised objective function value (as would be expected), and this also corresponds to a reduction in test domain model accuracy.
@@ Appendix D MMD analysis on Camelyon17
Split type Accuracy (%) MMD ×1000
@@
Linear k-means ($g\in\mathcal{Y\times D}$) 86.5 ± 1.3 97.0 ± 6.2
RBF kernel k-means ($g\in\mathcal{Y\times D}$) 85.6 ± 1.3 183.1 ± 2.0
Linear k-means ($g\in\mathcal{Y}$) 87.2 ± 0.2 108.3 ± 2.6
RBF kernel k-means ($g\in\mathcal{Y}$) 87.3 ± 0.7 206.0 ± 3.4

### v3

@@ 5 Experiments
The benefits of any new model selection method can only be verified when the oracle criterion suggests there is “room for improvement” over a basic random split, i.e., there is a performance gap between the two. Thus, the experiments described in this section are set up to reflect this scenario (the limitations of this are discussed further in Section 6). For example, it is noticed that UDA tends to exhibit a larger gap than DG, and this is especially pronounced (perhaps unsurprisingly) for adversarial algorithms, which tend to be more sensitive to hyperparameter choices.
Two batches of experiments are run, to reflect both the UDA and DG settings. Each batch comprises an identical training setup applied to 3 different datasets. For the DG experiments, models are trained using the CORAL algorithm (Sun & Saenko, 2016), with the clustering performed using constraints (4). For the UDA experiments, the DANN algorithm (Ganin et al., 2015) is used to adapt to an additional, unlabelled subset of test domain samples, as well as to align the training domains to each other. As DANN was observed to be more sensitive to domain imbalances, the validation split is set to preserve domain distributions, i.e., using constraints (5).
In this work, we fix the RBF bandwidth to $\gamma=1$. To assess the sensitivity of our results to the choice of kernel, we also test a linear kernel $\kappa(x,y)=x^{T}y$, which is equivalent to linear k-means clustering.
All feature extractors are finetuned on the entirety of $S$ before the features are computed for the clustering, regardless of pretraining. Experiments are conducted using the DomainBed framework (Gulrajani & Lopez-Paz, 2021). This means all-but-one of the domains are placed in the development set, and the remaining “evaluation” domain is randomly split into a UDA set (for adaptation, unused for the DG experiments), and an independent test set used to determine final accuracy values. Every domain is tested 3 times for reproducibility, each time with a different random seed for model initialisation, hyperparameter search and other stochastic variables. The reported accuracy values are averages over all domains and repeats. Further training and hyperparameter search details are given in Appendix C; unless otherwise stated, the remaining details all follow the Domainbed default options. The Gurobi Optimizer (Gurobi Optimization LLC, 2023) is used to solve the LPs. In total, the experiments involve training 5,160 models, requiring around 100 GPU-days of computation.
@@ 5.1 Datasets
The datasets represent a range of domain shift problems encompassing both image and audio classification tasks. In addition to the covariate shifts which occur across all datasets, the two ecological datasets (Humpbacks and TerraIncognita) also contain significant conditional shifts due to the open-set and sometimes annotator-dependent nature of wildlife monitoring data. With the exception of Camelyon17, the datasets are all small enough that the entire kernel matrix can be computed. So, Camelyon17 is the only dataset for which the Nyström method is applied.
Camelyon17-WILDS (Bándi et al., 2019; Koh et al., 2021) tumour detection in tissue samples across 5 hospitals, 2 classes and 455,954 samples. In keeping with the WILDS setup, the model in this case is trained from scratch rather than using pretrained weights. However, note that these results still cannot be compared directly with results from WILDS, as the DomainBed setup does not match exactly. License: CC0.
Humpbacks (Napoli & White, 2023) detection of humpback whale vocalisations across 4 recording locations, 2 classes and 43,385 samples. This is the only dataset not to use the ResNet-18 architecture, and instead uses a custom CNN architecture and acoustic front-end described in Appendix C. License: Proprietary.
SVIRO (Dias Da Cruz et al., 2020) classification of vehicle rear seat occupancy across 10 car models, 7 classes and a balanced subset of 24,500 samples of the original dataset. License: CC BY-NC-SA 4.0.
VLCS (Fang et al., 2013) object classification across 4 image datasets, 5 classes and 10,729 samples. License: unknown.
PACS (Li et al., 2017) object classification across 4 image styles (photos, art, cartoons, and sketches), 7 classes and 9,991 samples. License: unknown.
Terra Incognita (Beery et al., 2018) classification of wild animals across 4 camera trap locations, 10 classes and 24,788 samples. License: CDLA-Permissive 1.0.
@@ 5.2 Results
In total, 6 model selection methods are compared. These are: the random split; the leave-one-domain-out split; the test domain (oracle) validation set; a random split followed by mix-up augmentation on the validation set (Lu et al., 2023); and cluster-based splits using linear and kernel k-means clustering. The random and random-plus-mix-up splits are stratified by domain. The validation and test accuracies are class-balanced for all methods.
The results are shown in Table 1, along with standard errors (the standard errors are as computed by DomainBed, and capture variability in the overall experimental run, including random seeds and across domains). In the following sections, a 95% confidence level is used when verifying whether two values have a statistically significant difference, which corresponds to non-overlapping confidence intervals of 1.96 times the standard error (assuming normally distributed errors).
@@
Split type Camelyon Humpbacks SVIRO VLCS PACS TerraInc
@@
The range of possible performance improvement differs by dataset, as determined by the gap between the random split and oracle criterion; accuracy values should be considered relative to this scale when averaging across datasets. Therefore, a column of average normalised values is also shown, where each dataset is shifted and scaled such that the random split has a value of 0 and the oracle a value of 100.
@@
Where the validation set comprises multiple domains, model selection is based on average validation accuracy across these domains, as is the DomainBed default. Results based on worst-domain accuracy are also shown in Appendix D. In theory, worst-domain accuracy should provide an additional layer of distributional robustness with respect to the domain labels (Sagawa et al., 2019). However, no significant difference in test accuracy overall is observed if worst-case validation accuracy is used, although the standard error does increase. It is possible that the uncertainty set $\mathcal{Q}=\mathcal{D}^{S}$ has too few degrees of freedom for DRO to confer any meaningful robustness in this case.
@@
On average, the cluster-based splits provide a net absolute accuracy gain of around 3 percentage points compared to the random split, and 2 percentage points gain compared to leave-one-domain out validation. In relative terms, clustering is observed to close around 50% of the gap between the random split and oracle criterion, compared to 28% for leave-one-domain-out validation and 10% for mix-up. Overall, performance is slightly higher using the linear kernel than with the RBF kernel, although this is within margin of error. It is possible that the latter may be improved with more careful tuning of $\gamma$. It is noted that both clustering methods are also within margin of error of leave-one-domain-out validation.
5.3 Ablation study on VLCS
An ablation study conducted on the VLCS dataset is shown in Table 2. This shows the effects of finetuning the feature extractor before clustering, the Nyström approximation, and the different constraint sets (4) and (5). For this dataset, use of the Nyström approximation, as well as additional finetuning of $\theta_{F}$, are not observed to have significant effects on test accuracy. For the linear kernel, clustering with constraints (4) performs significantly lower than using constraints (5), however, for the RBF kernel, this difference is not observed.
@@ 5.3 Ablation study on VLCS
5.4 MMD analysis
As stated in Section 1, the motivation for cluster-based splits is the notion that increasing the MMD between the training and validation sets increases test domain accuracy. To provide empirical support for this, these two variables are plotted against each other in Figure 1, for each dataset. As for the clustering, an RBF kernel is used with $\gamma=1$. Again, the Nyström method is used to estimate the MMD for the Camelyon17 dataset due to its size. Each subplot in Figure 1 shows a different dataset, and each point in a subplot represents one of the 5 split types (not including the oracle), averaged across the domains and 3 repeats. Standard errors were sufficiently small (i.e., at least 2 order of magnitude smaller than the data range for all subplots) that error bars were not included. The correlation coefficients and associated $p$-values are also shown. As only monotonic associations are being tested for, Spearman’s rank correlation $\rho$ is used.
The correlation is clearest for Camelyon17, possibly because the dataset is large enough that a low-noise estimate of the MMD is possible. The results for this dataset are given in tabular form in Table 3, along with an additional ablation study comparing the different clustering constraints (4) and (5). It can be seen that the additional constraint on the clustering (i.e., taking $g\in\mathcal{Y\times D}$) reduces the optimised objective function value (as would be expected), and this also corresponds to a reduction in test domain model accuracy.
The evidence in this section supports the proposition that the validation split should be attempting to maximise the MMD between the training and validation sets, and that this is more effective than using the same metadata-based splitting rule as the test split, as the leave-one-domain-out split intends to do.
@@ 5.4 MMD analysis
Split type (Constraint) Accuracy (%) MMD ×1000
@@
Linear k-means (5) 86.5 ± 1.3 97.0 ± 6.2
RBF kernel k-means (5) 85.6 ± 1.3 183.1 ± 2.0
Linear k-means (4) 87.2 ± 0.2 108.3 ± 2.6
RBF kernel k-means (4) 87.3 ± 0.7 206.0 ± 3.4
5.5 Assessing the effect of the changing training distribution
As evidenced in (3), the use of non-random data splitting introduces a confounding variable to the experiments: as both the training and validation distributions are dependent on the split, it is possible the test accuracy is being influenced by the model training, as well as the validation. To support the claim that cluster-based splitting results in more generalisable model selection, it is necessary to decouple the effects of these interventions, and show that it is indeed the improved model validation, and not the training, that is improving test accuracy.
DG experiments UDA experiments
Split type Camelyon17 Humpbacks SVIRO VLCS PACS TerraInc Average
Leave-one-domain-out
85.4 ± 0.4 82.1 ± 1.3 99.0 ± 0.1 73.8 ± 0.3 82.2 ± 0.5 38.9 ± 1.2 76.9 ± 0.3
Linear k-means
88.0 ± 0.3 85.4 ± 0.8 99.1 ± 0.1 77.0 ± 0.3 85.7 ± 0.3 47.1 ± 0.9 80.4 ± 0.2
RBF kernel k-means
88.6 ± 0.6 85.0 ± 0.7 99.2 ± 0.1 77.9 ± 0.5 85.7 ± 0.3 44.5 ± 0.6 80.2 ± 0.2
These results show that performing a cluster-based split has no significant effect on the generalisation power of a model when the hyperparameters are being chosen via the oracle criterion. Therefore, it can be concluded that the performance improvements of cluster-based splitting seen in Table 1 come entirely from the model selection, and not from the training. On the other hand, test accuracy for the leave-one-domain-out split does significantly reduce. This finding may help to explain why metadata-based splits have been found to underperform random splits in some works (Gulrajani & Lopez-Paz, 2021): the increased robustness of OOD validation is not enough to counterbalance the reduced robustness of training on fewer domains.
5.6 Convergence behaviour of Algorithm 1
As mentioned previously, the proposed split method is similar in principle to the one of Wecker et al. (2020). One important implementational distinction is the manner by which the constraints are enforced. We have claimed that the greedy approach of Wecker et al. (2020) is susceptible to local minima and thus tends to converge to inferior solutions. To support this claim, this section presents an experiment to compare the quality of the optima attained by the two methods. We use the official implementation of Wecker et al. (2020), which uses linear k-means with label-only constraints (4), and configure our method similarly for fair comparison. For each domain in the VLCS dataset, we run both algorithms 20 times with different random initialisations, and $h=0.5$. The attained clustering objectives $\Psi(T,V)$ are shown in Figure 2. It can be seen that our method has significantly less spread than the one of Wecker et al. (2020) (around 1000 times less for Domain 3, 100 times less for Domains 1 and 4, and 1.8 times less for Domain 2), and the mean losses also tend to be lower. These findings suggest that the greedy method is indeed getting trapped in local minima.
Figure 2: The attained loss of our clustering algorithm vs that of Wecker et al. (2020), on the Humpbacks dataset, by domain.
5.7 Assessing the effect of the number of classes
As can be inferred from the problem formulation, a large number of classes may limit the effectiveness of the method. Specifically, with more classes, the partitioning problem is less flexible due to the increased number of clustering constraints, which reduces the upper bound on the MMD which can be achieved. For example, in an extreme case with only two examples per class, a random split (stratified by class) and a clustering-based split would be equivalent. Thus, this section presents an experiment to analyse the effect of the number of classes on MMD($T,V$). Specifically, using a Humpbacks subset of 8,000 examples, we generate synthetic class labels using another round of kernel k-means clustering, with a constraint to ensure equal class sizes. Then, we perform a kernel k-means data split using the synthetic labels, and the label-only constraints (4). Table 5 shows how the MMDs change as the number of classes varies between 1 and 500. We also report the MMDs for two of the baselines: the random split and the metadata-based split.
Table 5: The MMDs of the kernel k-means clustering data split by number of classes, and two baselines.
Number of classes/
Baselines
MMD($T,V$) MMD($T,E$) MMD($V,E$)
1 0.221 ± 0.016 0.152 ± 0.042 0.148 ± 0.040
2 0.187 ± 0.007 0.154 ± 0.048 0.129 ± 0.048
5 0.177 ± 0.007 0.121 ± 0.037 0.157 ± 0.034
10 0.161 ± 0.006 0.132 ± 0.020 0.138 ± 0.021
20 0.159 ± 0.005 0.132 ± 0.021 0.137 ± 0.019
50 0.145 ± 0.002 0.137 ± 0.031 0.125 ± 0.031
100 0.138 ± 0.004 0.119 ± 0.019 0.140 ± 0.017
500 0.129 ± 0.001 0.123 ± 0.027 0.130 ± 0.027
Random 0.001 ± 0.000 0.095 ± 0.064 0.095 ± 0.064
Leave-one-domain-out 0.052 ± 0.034 0.053 ± 0.036 0.135 ± 0.087
To determine a threshold for the maximum number of classes for which a clustering-based split is still valid, we can consider the range for which MMD($T,V$) is higher than MMD($T,E$) and MMD($V,E$), since this is the assumption made by DRO. We consider both MMD($T,E$) and MMD($V,E$) since the data split is symmetric (it does not matter which way around $T$ and $V$ are labelled). Note that MMD($T,E$) and MMD($V,E$) also tend to reduce with MMD($T,V$) due to the triangle inequality. In this case, it can be seen that this threshold occurs at around 100 classes. Although, it is noted that even for larger class numbers, the clustering-based split still yields significantly higher MMD($T,V$) than both baselines.
6 Limitations
As with all robust optimisation, a “no free lunch” theorem (Wolpert & Macready, 1997) applies to model selection as well: using a worst-case split loses the guarantee that the selected model will be optimal in the nominal (ID) case. If no domain shift is expected, an ID validation set is clearly preferable; in general, robust strategies may be inappropriate if the uncertainty is minimal, so the nature of the problem should be considered prior to choosing a methodology. Additionally, if unlabelled test domain data is available at split time, this should also be factored in when choosing the validation set (e.g., using the method of Napoli & White (2025)). However, this is not applicable if the hyperparameters need to be decided before the test domain is accessible, e.g., for test-time adaptation, and our experiments show that the worst-case split is still valid in the UDA case. The motivations and considerations for using worst-case optimisation are discussed thoroughly in Sagawa et al. (2019). It is also noted that in the nominal case, the favourable training distribution will be the dominating factor on test accuracy compared to the model selection, and so sub-optimality of the selection method will be less of an issue.
On evaluation, using $\gamma=1$ for the RBF kernel bandwidth was not an appropriate choice, especially in the high-dimensional case. A more practical option would be to use $\gamma=1/d_{F}$, where $d_{F}$ is the dimensionality of the clustering features (Liu et al., 2020).
As stated in Section 1, a major advantage of the clustering-based validation split is that it is not dependent on domain metadata. However, even with the domain constraint removed or with metadata-independent training algorithms (e.g., empirical risk minimisation), the specific experimental setup in this paper precludes the ability to test this method in truly metadata-free settings. Due to the fundamental design of the DomainBed framework, domain metadata can still be leveraged (both implicitly and explicitly) through several avenues. For example, models are still trained on domain-balanced minibatches of data, and validation accuracy is averaged over the validation domains. The datasets themselves may also be unrealistically domain-balanced to begin with. However, the authors believe that these influences are mild enough that the overall trends observed in these experiments can reasonably be expected to hold in metadata-free settings as well.
7 Conclusion
The algorithm is not parameter-free; future work could include a data-driven method for selecting these parameters, for example with an additional layer of meta-tuning. The algorithm can also trivially be extended to k-fold cross-validation by increasing the number of clusters. Finally, an investigation into the use of prior knowledge-informed feature extraction (see Appendix A) would be of interest.
8 Acknowledgements
This work was supported by grants from BAE Systems and the Engineering and Physical Sciences Research Council. The authors acknowledge the use of the IRIDIS High Performance Computing Facility, and associated support services at the University of Southampton, in the completion of this work.
Appendix A Prior Knowledge-Informed Feature Extraction
Clustering-based splits may be rendered more effective if the features used for clustering are informed by domain knowledge on the underlying causes of the domain shift. In this section, we discuss various applications where domain knowledge may be available and propose relevant features which could leverage this information.
In visual tasks involving stylistic shifts, style features (Matsuura & Harada, 2020) or CLIP-derived tokens can explicitly capture these differences. In histopathology, tumour detection is affected by staining variations across tissue slides, so colour-based features such as hue histograms or colour deconvolution can be used. In wildlife monitoring, shifts due to habitat or camera placement can be isolated by using segmentation to extract backgrounds. Visual changes such as lighting or colour tone shifts can be captured through metrics like brightness, contrast, and colour temperature. In cases where image resolution or fidelity varies, edge density, compression or downsampling artefacts, or image quality scores may provide useful indicators. Audio tasks can suffer from variations in recording equipment or conditions; features such as spectral descriptors, microphone gain, and SNR estimates can help characterise these differences. Or, for speech recognition, speaker differences can be captured using accent classifiers, phonetic parameters, or word rate.
Appendix B Derivation of equation (11)
Based on the law of total variance, a fundamental theorem in ANOVA is that the total sum of squares of a set of points can be decomposed into the sum of squares within each cluster and the sum of squares of each cluster centroid to the overall centroid (weighted by cluster size):
$\operatorname{SSq}(S_{X})=\operatorname{SSq}(T_{X})+\operatorname{SSq}(V_{X})+|T_{X}|\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{S}}\right\|_{\mathcal{H}}^{2}+|V_{X}|\left\|\mu_{\mathbb{P}_{V}}-\mu_{\mathbb{P}_{S}}\right\|_{\mathcal{H}}^{2}.$
Next, using the fact that
$\mu_{\mathbb{P}_{S}}=\frac{|T_{X}|\mu_{\mathbb{P}_{T}}+|V_{X}|\mu_{\mathbb{P}_{V}}}{|T_{X}|+|V_{X}|},$
we have
$\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{S}}\right\|_{\mathcal{H}}^{2}$ $=\left\|\mu_{\mathbb{P}_{T}}-\frac{|T_{X}|\mu_{\mathbb{P}_{T}}+|V_{X}|\mu_{\mathbb{P}_{V}}}{|T_{X}|+|V_{X}|}\right\|_{\mathcal{H}}^{2}$
$=\left\|\frac{\mu_{\mathbb{P}_{T}}(|T_{X}|+|V_{X}|)-|T_{X}|\mu_{\mathbb{P}_{T}}-|V_{X}|\mu_{\mathbb{P}_{V}}}{|T_{X}|+|V_{X}|}\right\|_{\mathcal{H}}^{2}$
$=\left\|\frac{|V_{X}|(\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{V}})}{|T_{X}|+|V_{X}|}\right\|_{\mathcal{H}}^{2}$
$=\frac{|V_{X}|^{2}}{(|T_{X}|+|V_{X}|)^{2}}\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{V}}\right\|_{\mathcal{H}}^{2}.$
Similarly,
$\left\|\mu_{\mathbb{P}_{V}}-\mu_{\mathbb{P}_{S}}\right\|_{\mathcal{H}}^{2}=\frac{|T_{X}|^{2}}{(|T_{X}|+|V_{X}|)^{2}}\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{V}}\right\|_{\mathcal{H}}^{2}.$
Substituting and simplifying:
$\operatorname{SSq}(S_{X})$ $=\operatorname{SSq}(T_{X})+\operatorname{SSq}(V_{X})+\frac{|V_{X}||T_{X}|^{2}+|T_{X}||V_{X}|^{2}}{(|T_{X}|+|V_{X}|)^{2}}\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{V}}\right\|_{\mathcal{H}}^{2}$
$=\operatorname{SSq}(T_{X})+\operatorname{SSq}(V_{X})+\frac{|T_{X}||V_{X}|}{|T_{X}|+|V_{X}|}\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{V}}\right\|_{\mathcal{H}}^{2}.$
Finally, rearranging:
$\left\|\mu_{\mathbb{P}_{T}}-\mu_{\mathbb{P}_{V}}\right\|_{\mathcal{H}}^{2}=\frac{|T_{X}|+|V_{X}|}{|T_{X}||V_{X}|}(\operatorname{SSq}(S_{X})-\operatorname{SSq}(T_{X})-\operatorname{SSq}(V_{X})),$
as required.
Appendix C Additional training and hyperparameter details
Experimental parameter Value
Hyperparameter random search size 10
Number of trials 3
Holdout fraction 0.2
UDA holdout fraction 0.5
Number of training steps 3000
Gaussian kernel bandwidth 1
Finetuning iterations before split 3000
Nyström subset size (if applicable) 2000
Architecture ResNet-18
Class balanced True
Table 7: Classification pipeline of the Humpbacks dataset. This follows the pipeline in (Napoli & White, 2023).
Step number Step detail
Acoustic front-end
1 Resample to 10 kHz
2 Mel-scale filter bank with 64 filters
3 Short-time Fourier transform with 100 ms FFT window, 50% overlap
4 Per-channel energy normalisation
5 Split into 3.92 s (128 pixel) analysis frames with 50% overlap
CNN
1 Conv2D (nodes=16, kernel=3x3, stride=2, activation=ReLU)
2 Conv2D (nodes=16, kernel=3x3, stride=2, activation=ReLU)
3 Conv2D (nodes=16, kernel=3x3, stride=2, activation=ReLU)
4 Conv2D (nodes=16, kernel=3x3, stride=2, activation=ReLU)
5 Global average-pooling 2D
6 Fully-connected layer
Appendix D Worst-case accuracy validation
DG experiments UDA experiments
Split type Camelyon Humpbacks SVIRO VLCS PACS TerraInc
Raw
average
Normalised
average
Leave-one-out
Linear k-means
RBF k-means